### v1

@@ 2.2 Broad Spectrum Applications
The rapid training capabilities of world models, exemplified by DayDreamer’s real-world robot learning efficiency, contrast starkly with traditional methods, highlighting the transformative potential of these models in accelerating learning processes and enhancing performance [42, 43].
Virtual scene and video generation emerge as pivotal applications, with SORA and Genie leading advancements in this space. SORA’s capability to produce coherent, high-definition videos from diverse prompts represents a significant step towards simulating complex world dynamics. Despite its challenges in physical interaction simulation, SORA’s consistent 3D spatial representation underscores its potential as a foundational world model [44]. Genie’s interactive environment generation, though not as advanced in video quality as SORA, introduces a novel dimension of user-driven world manipulation, offering a glimpse into future applications of world models in creating immersive, controllable virtual realities [45].
This comprehensive examination underscores the exceptional versatility and advancing frontiers of world models, illustrating their foundational role in driving innovation across gaming, robotics, virtual environment generation, and beyond. The convergence of these models’ capabilities with dynamic adaptation and multi-domain generalization heralds a new era of AI, where world models not only serve as tools for specific tasks but also as platforms for broader exploration, learning, and discovery.
{longtblr}
[ caption = Summary of Recent World Model Applications., , ] width = colspec = Q[46]Q[227]Q[188]Q[479], cell31 = r=2, cell51 = r=3, cell81 = r=2, cell101 = r=7, cell171 = r=6, cell231 = r=14, cell371 = r=5, hline1,42 = -0.08em, hline2-3,5,8,10,17,37 = -, Year & World Model Core Structure Contribution/Key Idea
2018 World models [13] MDN, RNN Pioneered the first world model for unsupervised learning of spatial and temporal representation of environments.
2019 PlaNet [24] RSSM Introduced a latent dynamics model featuring both deterministic and stochastic transition components for latent prediction.
DreamerV1 [23] RSSM Efficient policy learning through the propagation of analytic value gradients for optimized decision-making processes.
2020 DreamerV2 [25] RSSM Utilized discrete multi-class vectors instead of Gaussian distribution for more precise state representation predictions.
AWML [46] $\gamma$-Progress Developed a world model that employs curiosity-driven exploration.
MuZero [34] MCTS Integrated a tree-based search with a learned model for improved strategic planning and decision-making.
2021 EfficientZero [32] MCTS Crafted a sample efficient model-based visual RL algorithm built on MuZero.
Pathdreamer [40] RedNet, PatchGAN Engineered a world model for Indoor Vision-and-Language Navigation.
2022 MILE [47] Probabilistic Generative Model Adopted a model-based imitation learning approach for joint policy learning in autonomous driving.
SEM2 [48] RSSM Implemented filters to extract key features and balance data distribution, optimizing model performance and accuracy.
TRANSDREAMER [30] Transformer State-Space Model Introduced the first transformer-based stochastic world model.
IRIS [14] DAE, Autoregressive Transformer Constructed a world model combining a discrete autoencoder with an autoregressive transformer.
DreamingV2 [49] RSSM Introduced a novel approach by integrating both the discrete representation and the reconstruction-free objective.
DreamerPro [50] RSSM, Convolutional Encoder Enhanced resistance to visual interference, building upon the Dreamer model’s foundation.
MaxEnt Dreamer [51] RSSM, Maximum Entropy Implemented maximum entropy reinforcement learning to significantly enhance the world model’s exploration capabilities.
2023 DriveDreamer [52] ActionFormer, Diffusion Enabled controllable generation of driving scene videos equipped with simplified interactive features.
Adriver-I [53] MLLM, VDM Developed a world model that leverages image-text pairs, combined with large language models.
GAIA-1 [54] Autoregressive Transformer Crafted a system to generate realistic driving scenes using video, text, and action inputs, providing fine-grained control over the generated environment.
I-JEPA [29] JEPA Implemented a non-generative approach for self-supervised learning of images.
A-JEPA [28] JEPA Extended the I-JEPA self-supervised learning approach to the audio spectrum.
MC-JEPA [19] JEPA Employed self-supervised optical flow estimation from videos to jointly learn motion features.
2023 DreamerV3 [26] RSSM Adopted symlog predictions to seamlessly adapt across all ranges using a tailored set of hyperparameters.
Daydreamer [42] RSSM Utilized robots to deploy the Dreamer model in real-world settings, enabling direct learning without simulators.
Drive-WM [55] VAE, Diffusion Introduced the first multi-view, end-to-end autonomous driving world model.
TrafficBots [56] CVAE, Transformer Developed a world model for multimodal motion prediction and end-to-end driving.
Uniworld [57] Transformer Crafted a world model utilizing multi-view, label-free image-LiDAR pairs to model the surrounding scene.
Safe DreamerV3 [41] RSSM Introduced a world model specifically designed for safe reinforcement learning.
HarmonyDream [33] RSSM Incorporated a comprehensive approach by considering the loss of per-dimension scales, dimensions, and training dynamics together.
Dr.G [58] RSSM, DCL, RSID Employed a zero-shot MBRL framework that resists visual interference.
STORM [31] VAE, Transformer Combined transformers with variational autoencoders to introduce random noise into model-based reinforcement learning.
TWM [59] Transformer-XL Utilized the Transformer-XL architecture to learn long-term dependencies while staying computationally efficient.
DiffDreamer [60] Conditional Diffusion Model Developed a world model for scene extrapolation.
MUVO [61] MLP, GRU Utilized world model to predict 3D occupancy and camera and lidar observations.
SWIM [62] RSSM Trained a world model using videos of human hands to enable robots to interact with environment.
OccWorld [63] S-T Generative Transformer Generalized the world model to 3D Occupancy space.
2024 WorldDreamer [64] S-T Patchwise Transformer Facilitated learning of dynamic visual signal using attention.
S4WM [65] Parallelizable SSMs Modified world model backbones for improving long-term memory.
Genie [45] ST-Transformer Trained in an unsupervised manner on unlabeled Internet videos to generate interactive environments.
V-JEPA [66] JEPA Extent I-JEPA to videos for self-supervised video pretraining.
Think2Drive [67] RSSM Efficiently addressed CALAR v2 scenarios with world models.
3 World Models in Autonomous Driving
This section delves into the transformative application of world models within the autonomous driving sector, underscoring their pivotal contributions to environmental comprehension, dynamic prediction, and the elucidation of physical principles governing motion. As an emergent frontier in the application of world models, the autonomous driving domain presents unique challenges and opportunities for leveraging these advanced computational frameworks. Despite the burgeoning interest, the integration of world models into autonomous driving predominantly revolves around scenario generation, and planning and control mechanisms, areas ripe for exploration and innovation.
Figure 5: World Models in Autonomous Driving Pipelines.
3.1 Driving Scenario Generation
The acquisition of data in autonomous driving encounters substantial hurdles, including high costs associated with data collection and annotation, legal constraints, and safety considerations. World models, through self-supervised learning paradigms, offer a promising solution by enabling the extraction of valuable insights from vast quantities of unlabeled data, thereby enhancing model performance in a cost-effective manner. The application of world models in driving scenario generation is particularly noteworthy, as it facilitates the creation of varied and realistic driving environments. This capability significantly enriches training datasets, equipping autonomous systems with the robustness to navigate rare and intricate driving scenarios.
GAIA-1 [54] represents a novel autonomous generative AI model capable of creating realistic driving videos using video, text, and action inputs. Trained on extensive real-world driving data from British cities via Wayve, GAIA-1 learns and understands some real-world rules and key concepts in driving scenarios, including different types of vehicles, pedestrians, buildings, and infrastructure. It can predict and generate subsequent driving scenarios based on a few seconds of video input. Notably, the generated future driving scenarios are not closely tied to the prompt video but are based on GAIA-1’s understanding of world rules. Employing an autoregressive transformer network at its core, GAIA-1 predicts upcoming image tokens conditioned on the input image, text, and action tokens, then decodes these predictions back to pixel space. GAIA-1 can predict multiple potential futures and generate diverse videos or specific driving scenarios based on prompts (e.g., changing weather, scenes, traffic participants, vehicle actions), even including actions and scenes beyond its training set (e.g., forced entry onto sidewalks). This demonstrates its ability to understand and infer driving concepts not present in its training set. In the real world, such driving behaviors are hard to acquire data for due to their riskiness. Driving scenario generation allows for simulated testing, enriching data composition, enhancing system capabilities in complex scenarios, and better evaluating existing driving models. Moreover, GAIA-1 generates coherent actions and effectively captures the perspective influences of 3D geometric structures, showcasing its understanding of contextual information and physical rules.
DriveDreamer [52], also dedicated to driving scenario generation, differs from GAIA-1 as it is trained on the nuScenes dataset [68]. Its model inputs include elements like HD maps and 3D boxes, allowing for more precise control over driving scenario generation and deeper understanding, thus improving video generation quality. Additionally, DriveDreamer can generate future driving actions and corresponding predictive scenarios, aiding in decision-making.
ADriver-I employs current video frames and historical vision-action pairs as inputs for a multimodal large language model (MLLM) [69, 70] and a video latent diffusion model (VDM) [71]. The MLLM outputs control signals in an autoregressive manner, which serve as prompts for VDM to predict subsequent video outputs. Through continuous prediction cycles, ADriver-I achieves infinite driving in the predictive world.
Drawing inspiration from the success of large language models, WorldDreamer [64] approaches world modeling as an unsupervised visual sequence modeling challenge. It utilizes the STPT to concentrate attention on local patches within spatiotemporal windows. This focus facilitates dynamic learning of visual signals and accelerates the convergence of the training process. Although World Dreamer is a general-purpose video generation model, it has demonstrated exceptional performance in generating autonomous driving videos.
Beyond visual information, driving scenarios also include a plethora of critical physical data. MUVO [61] leverages the world model framework for the prediction and generation of driving scenes, integrates both LIDAR point clouds and visual inputs to predict videos, point clouds, and 3D occupancy grids of future driving scenes. Such a comprehensive approach significantly enhances the quality of predictions and generated outcomes. Particularly, the outcome 3D occupancy grids can be directly applied to downstream tasks. Taking it a step further, OccWorld [63] and Think2Drive [67] directly utilize the 3D occupancy information as system inputs to predict the evolution of the surrounding environment and plan the actions of the autonomous vehicles.
3.2 Planning and Control
Beyond scenario generation, world models are instrumental in learning within driving contexts, evaluating potential futures, and refining planning and control strategies. For instance, Model-based Imitation LEarning (MILE) [47] adopts a model-based imitation learning approach to jointly learn the dynamics model and driving behavior in CARLA from offline datasets. MILE employs a ’generalized inference algorithm’ for rational and visualizable imagination and prediction of future driving environments, using imagination to compensate for missing perceptual information. This ability enables planning future actions, allowing autonomous vehicles to operate without high-definition maps. In unexperienced test scenarios within the CARLA simulator, MILE significantly outperformed the state-of-the-art models, improving the Driving Score from 46 to 61 (compared to an expert data score of 88). MILE is characterized by long temporal and highly diversified future predictions. Using a decoder on the predicted future states, MILE demonstrates stable driving across various scenarios.
SEM2 [48], building upon RSSM, introduces the semantic masked world model to enhance the sampling efficiency and robustness of end-to-end autonomous driving. The authors contend that world models’ latent states contain too much task-irrelevant information, adversely affecting sampling efficiency and system robustness. Moreover, due to imbalanced training data, world models struggle to handle unexpected situations. To address these issues, a signature filter is introduced to extract key task features, reconstructing semantic masks using the filtered features. For data imbalance, a sampler is used to balance the data distribution. After trained and tested in CARLA, the performence of SEM2 showed substantial improvement over DreamerV2.
Considering that most autonomous vehicles typically have multiple cameras, multi-view modeling is also a crucial aspect for world models. Drive-WM [55] is the first multi-view world model designed to enhance the safety of end-to-end autonomous driving planning. Drive-WM, through multi-view and temporal modeling, jointly generates frames for multiple views, then predicts intermediate views from adjacent ones, significantly improving consistency across multiple views. Additionally, Drive-WM introduces a simple unified conditional interface, flexibly applying images, actions, text, and other conditions, simplifying the conditional generation process. Trained and validated on the nuScenes dataset [68] with six views, Drive-WM selects the best trajectory by sampling predicted candidate trajectories and using an image-based reward function. On the nuScenes dataset, using FID and FVD as standards, Drive-WM surpassed all other contemporary methods, demonstrating that predicting the future aids in planning for autonomous driving. Furthermore, consistent with GAIA-1, Drive-WM’s ability to navigate in non-drivable areas showcases the world model’s understanding and potential in handling out-of-domain cases. Besides, drawing inspiration from the seminal work of Alberto Elfes [72], UniWorld [57] introduces an innovative approach by utilizing multi-frame point cloud fusion as the ground truth for generating 4D occupancy labels. This method accounts for the temporal-spatial correlations present in images from multi-camera systems. By leveraging unlabeled image-lidar pairs, UniWorld undergoes pre-training for world models, significantly enhancing the understanding of environmental dynamics. When tested on the nuScenes dataset, UniWorld demonstrates notable improvements in IoU for tasks such as motion prediction and semantic scene completion compared to methods that rely on monocular pre-training.
TrafficBots [56], also an end-to-end autonomous driving model, places a greater emphasis on predicting the actions of individual agents within a scene. Conditioning on each agent’s destination, TrafficBots employs a Conditional Variational Autoencoder (CVAE) [73] to learn distinct personalities for each agent, thereby facilitating action prediction from a BEV perspective. Compared to alternative approaches, TrafficBots offers faster operation speeds and can scale to accommodate more agents. Although its performance may not yet rival state-of-the-art open-loop strategies, TrafficBots showcases the potential of closed-loop strategies for action prediction.
4 Challenges and Future Perspectives
4.1 Technical and Computational Challenges
1) Long-Term Scalable Memory Integration:
The quest to imbue world models with long-term, scalable memories that mirror the intricacies of human cognitive processes remains a formidable challenge in the realm of autonomous driving. The efficacy of these models is intrinsically linked to their architectural foundations, which currently face significant hurdles when tackling long-term tasks. These limitations obstruct the ability of models to retain and efficiently access information over prolonged periods, a critical capability for navigating the complex and dynamic environments encountered in autonomous driving. Contemporary models grapple with issues such as vanishing gradients [74] and catastrophic forgetting [75], which severely limit their long-term memory capabilities. Transformer architectures, despite their advancements in facilitating access to historical data through self-attention mechanisms, encounter obstacles in scalability and speed when processing lengthy sequences. Innovative approaches, exemplified by studies like TRANSDREAMER [30] and S4WM [65], explore alternative neural architectures aiming to surmount these barriers. Notably, S4WM has demonstrated superior performance in maintaining high-quality generation over sequences up to 500 steps, markedly surpassing traditional architectures. Yet, the decline in performance observed beyond 1000 steps accentuates the existing disparity between the capacities of artificial and biological memory systems.
2) Simulation-to-Real-World Generalization:
The disparity between simulated training environments and the multifaceted nature of real-world conditions presents a critical bottleneck in the evolution of autonomous driving technologies. Current simulation platforms, while advanced, fall short in perfectly mirroring the unpredictability and variability of real-world scenarios [76]. This discordance, manifesting in discrepancies in physical properties, sensor noise, and the occurrence of unforeseen events, critically undermines the applicability of world models trained solely in simulated environments.
The endeavor to develop world models capable of seamless generalization from simulation to real-world driving scenarios is paramount. This requires not only the refinement of simulation technologies to capture the subtleties and unpredictability of real-world environments more accurately but also the development of models that are inherently robust to the variances between simulated and real-world data. Enhancing the fidelity of simulations, employing domain adaptation techniques, and leveraging real-world data for continuous model refinement are potential pathways to achieving more effective generalization. Moreover, the integration of advanced sensory fusion techniques and the exploration of novel learning paradigms, such as meta-learning and reinforcement learning from diverse data sources, could further empower world models to adapt dynamically to the complexities of real-world driving. These advancements are pivotal for the realization of truly autonomous driving systems capable of navigating the myriad challenges posed by real-world environments with agility, accuracy, and safety.
4.2 Ethical and Safety Challenges
1) Decision-Making Accountability:
To achieve this, there is a pressing need for the integration of explainable AI (XAI) principles directly into the development of world models [77]. XAI aims to make AI decisions more interpretable to humans, providing clear, understandable explanations for the actions taken by autonomous vehicles. This involves not merely an exposition of the decision-making process but a comprehensive delineation of the ethical, logical, and practical considerations influencing these decisions. Implementing XAI within autonomous driving systems necessitates a multidisciplinary approach, drawing on expertise from AI development, ethics, legal standards, and user experience design [78]. Such collaborative efforts are essential to ensure that autonomous vehicles can engage in decision-making processes that are not only technically sound but also ethically defensible and socially acceptable.
2) Privacy and Data Integrity:
The reliance of autonomous driving technologies on extensive datasets for operation and continuous improvement brings to the forefront significant concerns regarding privacy and data security. The safeguarding of personal information against unauthorized access and breaches is a critical priority, requiring a robust framework for the ethical handling and protection of data.
Addressing these concerns involves a multifaceted strategy that extends beyond compliance with existing privacy regulations such as the General Data Protection Regulation (GDPR) in Europe [79]. It entails the establishment of stringent data governance policies that dictate the collection, processing, storage, and sharing of data. These policies should be designed to minimize data exposure and ensure data minimization principles, whereby only the data necessary for specific legitimate purposes is processed. Moreover, the deployment of advanced cybersecurity measures is critical to protect data integrity and confidentiality. This includes the utilization of encryption technologies, secure data storage solutions, and regular security audits to identify and mitigate potential vulnerabilities. Additionally, fostering transparency with users about how their data is collected, used, and protected is fundamental [80]. This can be achieved through clear, accessible privacy policies and mechanisms that empower users with control over their personal information, including options for data access, correction, and deletion [81].
4.3 Future Perspectives
1) Bridging Human Intuition and AI Precision
This integration enables autonomous vehicles to not only react to the immediate physical world but also to understand and adapt to the social and psychological dimensions of driving—interpreting gestures, predicting human behaviors, and making decisions that reflect a deeper understanding of human norms and expectations. For instance, a world model equipped with cognitive co-piloting capabilities could accurately predict pedestrian movements in urban settings, navigate social driving conventions at four-way stops, or adapt driving styles in response to passenger comfort and feedback.
2) Harmonizing Vehicles with the Urban Ecosystem
Another visionary perspective involves the role of world models in transforming autonomous vehicles into agents of ecological engineering, harmonizing with urban ecosystems through adaptive, responsive behaviors that contribute to environmental sustainability. World models, with their deep understanding of complex systems and dynamics, can enable autonomous vehicles to optimize routes and driving patterns not just for efficiency and safety, but also for environmental impact, such as minimizing emissions, reducing congestion, and promoting energy conservation.
Envision a scenario where world models enable a fleet of autonomous vehicles to dynamically adjust their operations based on real-time environmental conditions, traffic flow, and urban infrastructure status. These vehicles could collaboratively reroute to balance traffic loads across a city, reducing congestion and the urban heat island effect. They could prioritize routes and speeds that optimize fuel efficiency and reduce emissions, or even interface with smart city infrastructures to support energy grid balancing through vehicle-to-grid technologies.
5 Conclusion
Acknowledgement

### v2

@@ Acknowledgment
[Uncaptioned image] Yanchen Guan is currently a PhD student at the State Key Laboratory of Internet of Things for Smart City and the Department of Civil Engineering at the University of Macau. He holds a master’s degree in Mobility Engineering from Politecnico di Milano (2023) and a bachelor’s degree in Mechatronics Engineering from Harbin Institute of Technology (2019). His research primarily focuses on autonomous driving, intelligent transportation systems, mechanical structures, and data analysis.
[Uncaptioned image] Haicheng Liao (Student Member, IEEE) received the B.S. degree in software engineering from the University of Electronic Science and Technology of China (UESTC) in 2022. He is currently pursuing the Ph.D. degree at the State Key Laboratory of Internet of Things for Smart City and the Department of Computer and Information Science, University of Macau. His research interests include connected autonomous vehicles and the application of deep reinforcement learning to autonomous driving.
[Uncaptioned image] Zhenning Li (Member, IEEE) received his Ph.D. in Civil Engineering from the University of Hawaii at Manoa in 2019. Currently, he holds the position of Assistant Professor at the State Key Laboratory of Internet of Things for Smart City, as well as the Departments of Civil and Environmental Engineering and Computer and Information Science at the University of Macau, Macau SAR. He has published over 50 papers in top journals and conferences. His main areas of research focus on the intersection of connected autonomous vehicles and Big Data applications in urban transportation systems. He has been honored with several awards, including the TRB Best Young Researcher award and the CICTP Best Paper Award, amongst others.
[Uncaptioned image] Jia Hu (Member, IEEE) is currently working as a Zhongte Distinguished Chair of Cooperative Automation with the College of Transportation Engineering, Tongji University. Before joining Tongji University, he was a Research Associate with the Federal Highway Administration (FHWA), USA. He is an Editorial Board Member of the Journal of Intelligent Transportation Systems and the International Journal of Transportation Science and Technology. He is a member of TRB (a Division of the National Academies) Vehicle Highway Automation Commit-tee, the Freeway Operations Committee, the Simulation subcommittee of the Traffic Signal Systems Committee, and the Advanced Technologies Committee of the ASCE Transportation and Development Institute. He is the Chair of the Vehicle Automation and Connectivity Committee of the World Transport Convention. He is an Associate Editor of the American Society of Civil Engineers Journal of Transportation Engineering and IEEE OPEN JOURNAL OF INTELLIGENT TRANSPORTATION SYSTEMS.
[Uncaptioned image] Runze Yuan received a bachelor’s degree from Tsinghua University, Beijing, in 2017. He received the PhD degree from the University of Hawaii at Manoa, Honolulu, in 2023. He is currently a postdoc in the Department of Automation at Tsinghua University, Beijing. His main research areas include microscopic traffic flow modeling, artificial intelligence in transportation, traffic flow detection and control, and intelligent transportation systems.
[Uncaptioned image] Yunjian Li earned his Ph.D. in Applied Physics and Materials Engineering from the University of Macau in 2023. Presently, he serves as an Assistant Professor in the Faculty of Innovation Engineering at Macau University of Science and Technology. He has authored 15 SCI papers published in prestigious journals such as Nature Communications and Cement and Concrete Research.
[Uncaptioned image] Guohui Zhang is currently a Professor with the Department of Civil and Environmental Engineering, University of Hawai’i at Manoa. His primary research interests include transportation system resilience, transportation systems modeling, planning, and operation, traffic sensing and sensor data analytics, artificial intelligence in transportation, connected and autonomous vehicle systems, and transportation safety and security. He serves as a member of the American Society of Civil Engineers Connected and Autonomous Vehicle Impact Committee and the Transportation Research Board Information Systems and Technology Committee and as a panelist for multiple National Collaborative Highway Research Program projects.
[Uncaptioned image] Chengzhong Xu (Fellow, IEEE) received the Ph.D. degree from The University of Hong Kong, in 1993. He is currently the chair professor of computer science and the dean with the Faculty of Science and Technology, University of Macau. Prior to this, he was with the faculty at Wayne State University, USA, and the Shenzhen Institutes of Advanced Technology, Chinese Academy of Sciences, China. He has published more than 400 papers and more than 100 patents. His research interests include cloud computing and data-driven intelligent applications. He was the Best Paper awardee or the Nominee of ICPP2005, HPCA2013, HPDC2013, Cluster2015, GPC2018, UIC2018, and AIMS2019. He also won the Best Paper award of SoCC2021. He was the Chair of the IEEE Technical Committee on Distributed Processing from 2015 to 2019.